OpenStudy (anonymous):

Simplify 8 over the quantity of 2 plus 2i . –2 + 2i –2 – i –2 + i –2 – 2i

OpenStudy (jdoe0001):

and so... .checked yet what a "conjugate" is?

OpenStudy (jdoe0001):

and also, does \(\textit{difference of squares} \\ \quad \\ (a-b)(a+b) = a^2-b^2\qquad \qquad a^2-b^2 = (a-b)(a+b)\) ring a bell?